# Disentangled 4D Gaussian Splatting: Towards Faster and More Efficient Dynamic Scene Rendering

### v1

@@ Abstract
Novel-view synthesis (NVS) for dynamic scenes from 2D images presents significant challenges due to the spatial complexity and temporal variability of such scenes. Recently, inspired by the remarkable success of NVS using 3D Gaussian Splatting (3DGS), researchers have sought to extend 3D Gaussian models to four dimensions (4D) for dynamic novel-view synthesis. However, methods based on 4D rotation and scaling introduce spatiotemporal deformation into the 4D covariance matrix, necessitating the slicing of 4D Gaussians into 3D Gaussians. This process increases redundant computations as timestamps change—an inherent characteristic of dynamic scene rendering. Additionally, performing calculations on a four-dimensional matrix is computationally intensive. In this paper, we introduce Disentangled 4D Gaussian Splatting (Disentangled4DGS), a novel representation and rendering approach that disentangles temporal and spatial deformations, thereby eliminating the reliance on 4D matrix computations. We extend the 3DGS rendering process to 4D, enabling the projection of temporal and spatial deformations into dynamic 2D Gaussians in ray space. Consequently, our method facilitates faster dynamic scene synthesis. Moreover, it reduces storage requirements by at least 4.5% due to our efficient presentation method. Our approach achieves an unprecedented average rendering speed of 343 FPS at a resolution of $1352\times 1014$ on an RTX 3090 GPU, with experiments across multiple benchmarks demonstrating its competitive performance in both monocular and multi-view scenarios.
@@
Figure 0: We present Disentangled 4D Gaussian Splatting, a novel approach that takes videos as input and synthesizes dynamic novel-view images with fine details. The proposed method surpasses previous approaches in both rendering quality and speed, achieving an inference speed of 343 FPS on an RTX 3090 in the Plenoptic Dataset[14].
1 Introduction
(a) The slicing first 4D Gaussian Splatting
(b) Disentangled 4D Gaussian Splatting
Figure 1: Comparison between Slicing first 4D Gaussian Splatting and our Disentangled 4D Gaussian Splatting. The upper one is the slicing first 4D Gaussian Splatting method, which need to slicing the 4D Gaussian into 3D Gaussian, not only do they need to calculate four-dimensional matrices, but they also have to repeat slicing and projection operations. In contrast, our method preserves temporal information during the rendering process while more conveniently obtaining continuous images in the temporal dimension.
Reconstructing dynamic scenes from 2D images and synthesizing photo-realistic novel views in real-time, has been a long-standing goal in computer vision and graphics. This task has attracted increasing attention from both academia and industry because of its potential value in various applications, including film, gaming, and VR/AR. Recently, such as NeRF [17], have achieved photo-realistic static scene rendering. However NeRF based methods typically can not support real-time rendering even for static scenes, due to their dense sampling of millions of rays. Recently, 3D Gaussian Splatting (3DGS) [10] has attracted widespread attention because of its real-time rendering ability for photo-realistic images in static scene. By modeling three-dimensional scene with 3D Gaussian ellipsoids explicitly and employing fast rasterization technique, it avoids dense ray sampling and achieves true real-time rendering in static scene.
@@ 1 Introduction
Subsequently, many efforts [28, 26, 6, 28] have attempted to apply 3DGS to dynamic scenes. All these methods based on 3DGS can be categorized into two groups. The first group [28, 26] employs deep learning to implicitly model temporal changes in 3D Gaussian ellipsoids. To synthesize novel views of dynamic scenes, the deep learning model predicts modifications to the base 3D Gaussian ellipsoids, which are then adjusted to reconstruct the scene at a given timestamp. This updated scene is subsequently processed through the 3DGS rendering pipeline [10], where 3D Gaussian ellipsoids are projected into 2D Gaussian ellipses in ray space and efficiently rasterized to produce the final novel view image. However, since each Gaussian ellipsoid requires querying a deep learning model, these methods often struggle to achieve real-time rendering speeds for complex scenes. The second group [6, 28] extends 3D Gaussian ellipsoids into 4D Gaussian hyperspheres, modeling dynamic scenes using four-dimensional transformations, including rotation, scaling, and mean. As Fig.1(a) shown, to generate novel views at specific timestamps, these methods first slice the 4D Gaussian hypersphere into a 3D Gaussian ellipsoid corresponding to the given timestamp before proceeding with the 3DGS rendering pipeline. Despite their differences, both groups require transforming dynamic scenes into static representations before rendering. While this method suffices for static novel view synthesis, it is inadequate for dynamic scenes where timestamps change more frequently than viewpoints. In deep learning-based methods, temporal deformations are modeled implicitly, making direct projective transformation infeasible. Similarly, prior rotation and scaling based 4D Gaussian approaches embed temporal transformations within the four-dimensional covariance matrix, making direct projection computationally expensive and complex. Consequently, these methods must recompute the entire rendering process whenever the timestamp changes, leading to significant redundant calculations.
To reduce redundant computations, we propose Disentangled 4D Gaussian Splatting (Disentangled4DGS), an approach that differs from traditional rotation and scale based representation methods. Specifically, we represent 4D Gaussians using a combination of 3D Gaussian parameters, time scaling, and the velocity of the mean. Our method effectively disentangles the time-dependent components of the four-dimensional Gaussian sphere from other parameters, enabling projection into ray space without extensive and complex calculations. Compared to traditional methods that require incorporating a four-dimensional covariance matrix during slicing to obtain a three-dimensional Gaussian sphere for projection and rasterization as Fig.1(a) shown, operates directly on the projected 2D Gaussian when the timestamp changes as Fig.1(b) shown, thus avoiding the repetition of the process from 4D Gaussian to 2D Gaussian. Additionally, because our method requires computing only the covariance matrix of the three-dimensional Gaussian sphere rather than a four-dimensional matrix during rendering, it offers substantial computational speed advantages. Furthermore, since our representation method utilizes fewer numerical values than traditional approaches, it naturally reduces storage requirements by at least 4.5% compared to the rotation-plus-scaling method. We also provide a formal proof of equivalence between the two representations in the supplementary material. Therefore, prior scene representations can be transformed into our format, enabling faster rendering using our optimized process.
The nature of Gaussian functions can lead to artifacts when expressing high-frequency information [30]. However, high-frequency features are prevalent in both spatial (e.g. sharp objects) and temporal (e.g. sudden appearance or disappearance) dimensions. And to alleviate this artifact problem that may arise from spatial high-frequency parts, we adopted an edge loss based on Laplacian convolution kernel [25] which is widely utilized to extract high-frequency parts. Experimental results indicate that our loss function significantly resolves artifact issues, enhancing rendering quality.
In complex scenarios, frequency distribution differences between time and space domains often render a unified density control strategy unsuitable. To address this, we decouple the traditional density control into temporal splitting strategy and spatial domain splitting strategy. Our approach delivers improved performance in dynamic scenes, as demonstrated by experimental results.
@@
  • We propose a disentangled representation for 4D Gaussians that separates time-dependent components from other parameters, eliminating the need for four-dimensional matrix operations. This approach reduces storage consumption by 4.5% compared to traditional methods while maintaining compatibility with existing representations through transformation.
@@
  • We propose an edge loss based on a Laplacian convolution kernel and a temporal splitting strategy to reduce visual artifacts.
@@
  • Disentangled4DGS achieves unprecedented rendering performance, generating $1353\times 1014$ resolution videos at an average of 343 FPS on an RTX 3090 in Plenoptic Video Dataset [14], as shown in Fig.1. Both quantitative and qualitative evaluations demonstrate its superiority over previous methods, setting new benchmarks in rendering quality and speed.
2 Related work
2.1 Novel view synthesis for static scenes
@@
In recent years, novel view synthesis has received widespread attention. Prior work formalized concepts like lumigraph [3] or light-field [12] and generating novel-view images through interpolating the existing views. Although traditional methods are efficient, they require densely captured images in complex scenes. Ben et al. [18] initiated a significant trend by utilizing a Multi-Layer Perceptron (MLP) to learn the radiance field and applying volumetric rendering for photo-realistic image synthesis from any viewpoint. Subsequent efforts have aimed to accelerate training and rendering [5, 29, 19] and enhance rendering quality by addressing existing issues in the NeRF [18], such as aliasing and reflection. However, NeRF [18] based methods involve querying the MLP for hundreds of points per ray, significantly consuming time. In contrast, Kerbl et al.’s 3D Gaussian Splatting (3DGS) [10] offers a novel framework enabling real-time, high-fidelity novel view synthesis for complex scenes. Recent advancements [30] have focused on enhancing both rendering speed and the quality of synthesized images.
2.2 Novel view synthesis for dynamic scenes
@@
Unlike static scenes, novel view synthesis for dynamic scenes necessitates accurate modeling of geometric shapes and colors, as well as their temporal changes. Inspired by NeRF’s success with static scenes, some works [19, 20] have attempted to extend NeRF [18] for dynamic environments, focusing on improving volume and spatiotemporal encoding for efficient dynamic scene modeling. The real-time rendering capabilities of 3DGS for photo-realistic images have drawn widespread attention. Some research [26] focuses on modeling temporal changes in 3DGS [10] using deep learning, yielding high-quality novel view synthesis. However, since each Gaussian sphere necessitates querying the model to capture temporal changes, the rendering speed is slower than the original 3DGS.
2.3 Dynamic 3D Gaussians
Figure 2: Rendering pipeline of our Disentangled 4DGS. After initialization, we first project the 3D Gaussians and the velocity of mean orthogonally to the timeline, obtaining a 2D Gaussian sphere with velocity in ray space. Then the projected 2D Gaussians with velocity are sliced to obtain the static 2D Gaussian in ray space and utilize rasterization to get the image in the end. The gradients from loss are back-propagated to optimize the 4D Gaussians and guide the adaptive density control
@@ 2.3 Dynamic 3D Gaussians
There has been significant progress in extending 3DGS to dynamic scenes [6, 28]. Yang et al. [28] characterize 4D Gaussian hyperspheres by expanding the quaternions and 3D scaling vectors of 3D Gaussian spheres into dual quaternions and 4D scaling vectors, and they innovatively use four-dimensional spherical harmonics to capture temporal and spatial color variations. Duan et al. [6] employ a spatiotemporal decoupled 16-dimensional vector instead of dual quaternions to represent rotation in four-dimensional space, reducing degrees of freedom and optimization complexity. These methods render images faster than deep learning-based approaches. But all of these methods need to get a static scenes at a timestamp represented by 3DGS firstly then follow the rendering pipeline of 3DGS to get a image of novel view at the timestamp. Thus these methods have low efficiency when timestamp changes frequently, which often occurs in dynamic scenarios. In contrast, by projecting the Gaussian function in ray space firstly, we can delay the time dimension in rendering, thereby reducing duplicate calculations when time changes.
3 Method
@@
3.1 Preliminary of 3D Gaussian Splatting
@@
To ensure that the covariance matrix $\Sigma$ is positive definite and symmetric. 3DGS optimizes the scaling matrix $\mathbf{S}$ and rotation matrix $\mathbf{R}$ to replace direct optimizing covariance matrix. The covariance matrix $\Sigma$ can be formulated by $\mathbf{S}$ and $\mathbf{R}$ as:
$\Sigma=\mathbf{RS}\mathbf{S}^{T}\mathbf{R}^{T}$ (2)
where $\mathbf{S}=diag(s_{x},s_{y},s_{z})\in\mathbb{R}^{3}$ and $\mathbf{R}\in SO(3)$ are stored as a 3D vector and quaternion. To represent the color of each Gaussian ellipsoid, 3DGS employs two sets of additional learnable parameters including opacity $o\in(0,1)$ and spherical harmonic (SH) coefficients in $\mathbb{R}^{k}$ representing the color changed by views, where $k$ is related to SH order. The initial process of 3DGS rendering is computing the covariance matrix $\Sigma$. A view matrix then is used to transform the Gaussian ellipsoids from world space to camera space. Because this transformation is an affine transformation, the ellipsoids are still Gaussian ellipsoids. To get the 2D reconstruction kernel in ray space [32], we need to perform a mapping $x=\phi(\mathbf{P_{3D}})$ on Gaussian ellipsoids, which can be formulated as:
@@
where $\mathbf{J}_{k}$ is the Jacobian matrix of $\phi$ at the point $\mathbf{t}_{k}$, which can be described as:
$\mathbf{J}_{k}=\frac{\partial\phi}{\partial\mathbf{t}}(\mathbf{P}_{k})=\begin{pmatrix}1/P_{k,2}&0&-P_{k,0}/P^{2}_{k,2}\\ 0&1/P_{k,2}&-k_{k,1}/P^{2}_{k,2}\\ P_{k,0}/l^{{}^{\prime}}&P_{k,1}/l^{{}^{\prime}}&P_{k,2}/l^{{}^{\prime}}\end{pmatrix}$ (5)
where $l^{{}^{\prime}}=\left\|(P_{k,0},P_{k,1},P_{k,2})^{T}\right\|$. In the end, the color of each pixel in the image is calculated by blending Gaussians sorted by their depths:
@@
3.2 Disentangled 4D Gaussian Splatting
We now discuss our Disentangled 4D Gaussian Splatting (Disentangled4DGS) algorithm, as illustrated in Fig. 2. Specifically, we model the 4D Gaussian using 3D Gaussian (Sec. 3.2.1), time scaling and velocity of mean, and project the 4D Gaussian into ray space [32] (Sec. 3.2.2). We then slice the Gaussian in ray space and apply the fast rasterization technique to generate image.
3.2.1 Representation of 4D Gaussian
@@
$\Sigma_{4D}=\mathbf{R}_{4D}\mathbf{S}_{4D}\mathbf{S}^{T}_{4D}\mathbf{R}^{T}_{4D}=\begin{pmatrix}\mathbf{U}&\mathbf{V}\\ \mathbf{V^{T}}&\mathbf{W}\end{pmatrix}.$ (7)
Extending the 3D scaling matrix to 4D is straightforward; however, extending quaternion-based 3D rotation to 4D is more complex. Yang et al. [28] employs the dual quaternion and Duan et al. [6] adopt the 4D rotor to characterize the 4D rotations. Both of their slicing methods are based on conditional probability derivations, formulated as:
@@
As shown in Eq. 9 , $\mu_{3D}$ is influenced by the 4D covariance matrix $\Sigma_{4D}$ along the time dimension.Directly projecting the coupled scale $\mathbf{S}_{4D}$ and rotation $\mathbf{R}_{4D}$ is both difficult and computationally expensive. To address this, we decompose Eq. 9 into:
  • velocity of mean $\mathcal{V}_{3D}=\frac{\mathbf{V}}{\mathbf{W}}$
  • 3D Gaussian representation, where $\mathbf{U}-\frac{\mathbf{V}}{\mathbf{V}^{T}}$ serves as covariance matrix $\Sigma_{3D}$, and $(\mu_{x},\mu_{y},\mu_{z})^{T}$ serves as the mean vector $\mu_{3d}$
  • Temporal scaling, $S_{t}=\sqrt{\mathbf{W}}$ to ensure the variance remains positive
A detailed proof demonstrating the equivalence of these two representations is provided in the supplementary material. Using our representation, only the velocity of mean and temporal scaling are affected by deformations along the time dimension. Due to the projection direction is perpendicular to the time dimension to ensure that the length of time dimension is not affected. The temporal scale $S_{t}$ is unaffected by projection, while the velocity of mean $\mathcal{V}_{3D}$, can be easily projected similarly to the mean vector $\mu_{3d}$.
In summary, we presentation 4D Gaussian with:
@@
where $\mathbf{s}_{4D}$ is the scaling vector, $q$ represents the quaternion, and $\mathcal{V}_{3D}$ denotes the velocity of mean. Our presentation effectively disentangles temporal deformations, enabling for efficient and fast projection into the ray space [32]. Compared to traditional methods requiring 16 floating-point numbers, our representation only requires 15, thereby improving storage efficiency. Consequently, our approach reduces storage requirements by at least 4.5%.
3.2.2 The Differentiable Projection First Splatting for 4D Gaussian
Following the pipeline of 3D Gaussian Splatting (3DGS), the mean of a 4D Gaussian, denoted as $\mu_{4d}=(\mu_{3d},t)$, is tranformed from world space to camera space before projection, which can be expressed as:
$P_{3D}=\varphi(\mu_{3D}),P_{t}=t,$ (11)

### v3

# Disentangled 4D Gaussian Splatting: Rendering High-Resolution Dynamic World at 343 FPS

@@ Abstract
While dynamic novel view synthesis from 2D videos has seen progress, achieving efficient reconstruction and rendering of dynamic scenes remains a challenging task. In this paper, we introduce Disentangled 4D Gaussian Splatting (Disentangled4DGS), a novel representation and rendering pipeline that achieves real-time performance without compromising visual fidelity. Disentangled4DGS decouples the temporal and spatial components of 4D Gaussians, avoiding the need for slicing first and four-dimensional matrix calculations in prior methods. By projecting temporal and spatial deformations into dynamic 2D Gaussians and deferring temporal processing, we minimize redundant computations of 4DGS. Our approach also features a gradient-guided flow loss and temporal splitting strategy to reduce artifacts. Experiments demonstrate a significant improvement in rendering speed and quality, achieving 343 FPS when render $1352\times 1014$ resolution images on a single RTX 3090 while reducing storage requirements by at least 4.5%. Our approach sets a new benchmark for dynamic novel view synthesis, outperforming existing methods on both multi-view and monocular dynamic scene datasets.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@
Figure 0: We present Disentangled 4D Gaussian Splatting, a highly-efficient approach that renders 1352 × 1014 resolution images at 343 FPS on an RTX 3090 in the Plenoptic Dataset[2], surpassing previous approaches in both rendering quality and speed. Note that the x-axis is logarithmic scale.
I Introduction
Reconstructing dynamic scenes from 2D images and synthesizing photo-realistic novel views in real-time, has been a long-standing goal in computer vision and graphics. This task has attracted increasing attention from both academia and industry because of its potential value in various applications, including film, gaming, and VR/AR[3]. While promising progress has been made, achieving both high-quality and highly efficient rendering for dynamic scenes remains a substantial challenge. Building upon the success of 3DGS in static settings, researchers[1],[4],[5] have begun to explore its generalization to the spatio-temporal (4D) domain, aiming to achieve real-time rendering for dynamic scenes.
@@ I Introduction
However, existing 4DGS methods face significant efficiency limitations. Existing works on 4DGS can be categorized into two branches. The first branch [1],[4] employs deep learning to implicitly model temporal changes in 3D Gaussian ellipsoids. These methods predict variations to the base 3D Gaussian ellipsoids and synthesize novel views by projecting them into 2D Gaussian ellipses in ray space[6]. However, per-Gaussian network queries can be computationally expensive and become a bottleneck for real-time performance in large scenes. The second branch [5],[1] extends 3D Gaussian ellipsoids into 4D Gaussian hyperspheres, modeling dynamic scenes through four-dimensional transformations of rotation, scaling, and mean. As shown in Fig.1(a), to generate novel views at specific timestamps, these methods first temporally slice the 4D Gaussian hypersphere into a 3D Gaussian ellipsoid corresponding to the given timestamp before proceeding with the 3DGS rendering pipeline. Despite of their differences in design, a key limitation shared by both approaches is that they must process (or generate) a static 3D scene representation for each specific timestamp before the main rendering steps. Specifically, in the first branch, the implicit deep learning model does not support direct dynamic projection. In the second branch, temporal transformations embedded within the 4D covariance matrix make direct projection computationally expensive and complex[6]. Consequently, these methods must repeat the entire rendering process whenever the timestamp changes, leading to significant redundant computations.
To address the inefficiency caused by redundant computations in existing 4DGS methods[5],[1], we propose Disentangled 4D Gaussian Splatting (Disentangled4DGS). Our core innovation lies in a novel 4D Gaussian representation that disentangles time-dependent properties from other parameters, such as 3D Gaussian parameters, time scaling, and the velocity of the mean. This disentangled representation unlocks a more efficient rendering pipeline. Unlike existing ”slicing-first” methods (Fig. 1(a)) that rely on slicing the 4D Gaussian at each timestamp, bring redundant matrix operations with expensive computation, our method (Fig. 1(b)) allows for timestamp updates to directly affect the projected 2D Gaussians, eliminating the need to repeatedly execute the entire 4D-to-2D pipeline repeatedly. Benefiting from our uniquely designed Disentangled 4DGS, our method is memory-efficient and achieves substantial computational speedups by deferring temporal processing and avoiding complex 4D matrix operations.
Another challenge of 4DGS lies in accurately modeling object motions in complex dynamic scenes. In practice, object trajectories may become distorted, especially near motion boundaries, due to the lack of explicit geometric constraints on dynamic regions. To address this issue, we introduce a flow-gradient guided consistency loss that leverages image gradients as structural cues to regulate motion boundaries. By aligning the gradients of motion fields with image edges, this loss encourages that motion discontinuities occur only at true object boundaries while remaining smooth within homogeneous regions. Our experiments demonstrate that this constraint leads to more coherent object trajectories and improves rendering quality in dynamic scenes.
Furthermore, existing 4DGS approaches[7, 5, 1] often approximate object motions with first-order or second-order models. While such simplifications suffice for relatively smooth trajectories, they struggle with complex and non-linear dynamics such as sudden appearance, disappearance, or occlusion changes. To overcome this limitation, we introduce temporal splitting strategies that explicitly decouple complex motions into finer temporal segments, enabling a more accurate representation of non-linear object trajectories. Experimental results confirm that combining flow-gradient guided consistency with temporal splitting significantly enhances motion fidelity and overall rendering quality in dynamic environments.
@@
  • We propose a disentangled representation for 4D Gaussians that separates time-dependent components from other parameters, largely improving the efficiency of 4DGS.
@@
  • We propose a flow-gradient guided consistency loss and a temporal splitting strategy to better constrain object motions and improve rendering quality in dynamic scenes.
@@
  • Disentangled4DGS achieves unprecedented rendering performance, generating $1353\times 1014$ resolution videos at an average of 343 FPS on an RTX 3090 in Plenoptic Video Dataset [2], as shown in Fig.1. Both quantitative and qualitative evaluations demonstrate its superiority over previous methods, setting new benchmarks in rendering quality and efficiency.
(a) The ”slicing-first” 4D Gaussian Splatting
(b) Our disentangled 4D Gaussian Splatting
Figure 1: Comparison between ”slicing-first” 4D Gaussian Splatting and our Disentangled 4D Gaussian Splatting. The upper one is the slicing first 4D Gaussian Splatting method, which need to slice the 4D Gaussian into 3D Gaussian. This approach requires computing high-dimensional covariance matrices and performing repeated slicing and projection operations, leading to inefficiency and temporal discontinuity. In contrast, our ”projection-first” disentangled formulation preserves temporal information throughout the rendering pipeline, enabling efficient rasterization and continuous, temporally coherent image synthesis.
II Related work
II-A Novel view synthesis for static scenes
@@
In recent years, novel view synthesis has received widespread attention. Prior work formalized concepts like lumigraph [8] or light-field [9] and generating novel-view images through interpolating the existing views. Although traditional methods are efficient, they require densely captured images in complex scenes. Ben et al. [10] initiated a significant trend by utilizing a Multi-Layer Perceptron (MLP) to learn the radiance field and applying volumetric rendering for photo-realistic image synthesis from any viewpoint. Subsequent efforts have aimed to accelerate training and rendering [11],[12],[13],[14],[15],[16] and enhance rendering quality by addressing existing issues in the NeRF [10], such as aliasing and reflection. However, NeRF [10] based methods involve querying the MLP for hundreds of points per ray, significantly consuming time. In contrast, Kerbl et al.’s 3D Gaussian Splatting (3DGS) [17] offers a novel framework enabling real-time, high-fidelity novel view synthesis for complex scenes. Recent advancements [18],[19],[20],[21],[22],[23],[24] have focused on enhancing both rendering speed, sparse inputs and the quality of synthesized images.
II-B Novel view synthesis for dynamic scenes
@@
Unlike static scenes, novel view synthesis for dynamic scenes necessitates accurate modeling of geometric shapes and colors, as well as their temporal changes[25],[26],[27],[28]. Inspired by NeRF’s success with static scenes, some works [14],[29] have attempted to extend NeRF [10] for dynamic environments, focusing on improving volume and spatiotemporal encoding for efficient dynamic scene modeling. The real-time rendering capabilities of 3DGS for photo-realistic images have drawn widespread attention. Some research [4],[30],[31],[32],[33],[34],[35] focuses on modeling temporal changes in 3DGS [17] using deep learning, yielding high-quality novel view synthesis. However, since each Gaussian sphere necessitates querying the model to capture temporal changes, the rendering speed is slower than the original 3DGS.
II-C Dynamic 3D Gaussians
@@
There has been significant progress in extending 3DGS to dynamic scenes [5],[1]. Yang et al. [1] characterize 4D Gaussian hyperspheres by expanding the quaternions and 3D scaling vectors of 3D Gaussian spheres into dual quaternions and 4D scaling vectors, and they innovatively use four-dimensional spherical harmonics to capture temporal and spatial color variations. Duan et al. [5] employ a spatiotemporal decoupled 16-dimensional vector to represent rotation in four-dimensional space, reducing degrees of freedom and optimization complexity. These methods render images faster than deep learning-based approaches. But all of these methods need to get a static scene for each timestamp before following the 3DGS rendering pipeline. Thus these methods have low efficiency when timestamp changes frequently, which often occurs in dynamic scenarios. In contrast, by projecting the Gaussian function in ray space firstly, we can delay the time dimension in rendering, thereby reducing duplicate calculations when time changes.
III Method
@@
III-A Preliminary of 3D Gaussian Splatting
@@
The initial process of 3DGS rendering is computing the covariance matrix $\Sigma$ through scale $s_{3D}$ and quaternion $q$. A view matrix then is used to transform the Gaussian ellipsoids from world space to camera space. Because this transformation is an affine transformation, the ellipsoids are still Gaussian ellipsoids. To get the 2D reconstruction kernel in ray space [6], we need to perform a mapping $x=\phi(\mathbf{P_{3D}})$ on Gaussian ellipsoids, which can be formulated as:
@@
where $\mathbf{J}_{k}$ is the Jacobian matrix of $\phi$ at the point $\mathbf{t}_{k}$. In the end, the color of each pixel in the image is calculated by blending Gaussians sorted by their depths:
@@
Figure 2: Rendering pipeline of our Disentangled 4DGS. After initialization, we first project the 3D Gaussians and the velocity of mean orthogonally to the timeline, obtaining a 2D Gaussian sphere with velocity in ray space. Then the projected 2D Gaussians with velocity are sliced to obtain the static 2D Gaussian in ray space and utilize rasterization to produce the image. The gradients from loss are back-propagated to optimize the 4D Gaussians and guide the adaptive density control.
III-B Disentangled 4D Gaussian Splatting
We now introduce our Disentangled 4D Gaussian Splatting (Disentangled4DGS) algorithm, as illustrated in Fig. 2. Specifically, we model the 4D Gaussian using 3D Gaussian (Sec. 3.2.1), time scaling, and velocity of mean. The presentation is then projected into ray space [6] (Sec. 3.2.2). We then slice the Gaussian in ray space and apply the fast rasterization technique to render images.
III-B1 Representation of 4D Gaussian
@@
$\Sigma_{4D}=\mathbf{R}_{4D}\mathbf{S}_{4D}\mathbf{S}^{T}_{4D}\mathbf{R}^{T}_{4D}=\begin{pmatrix}\mathbf{U}&\mathbf{V}\\ \mathbf{V^{T}}&\mathbf{W}\end{pmatrix}$, (5)
where $\mathbf{U}$ is the covariance matrix of base 3D Gaussian, $\mathbf{V}$ is the Spatiotemporal covariance, $\mathbf{W}$ is the temporal variance. Extending the 3D scaling matrix to 4D is straightforward; however, extending quaternion-based 3D rotation to 4D is more complex due to the intrinsic coupling between spatial and temporal rotations in the 4D domain. Yang et al. [1] employs the dual quaternion and Duan et al. [5] adopt the 4D rotor to characterize the 4D rotations. Both of their slicing methods are based on conditional probability derivations, formulated as:
@@
As shown in Eq. 7, the resulting 3D mean $\mu_{3D}$ is dependent on the components of the 4D covariance matrix $\Sigma_{4D}$ along the time dimension, even though the 3D mean itself is defined only in spatial coordinates. This inherent spatial-temporal coupling made the direct projection of the combined scale $\mathbf{s}_{4D}$ and rotation $\mathbf{R}_{4D}$ both challenging and computationally expensive. Besides, the projection of 4D covariance matrix $\Sigma_{4D}$ is not an affine transformation, thus a complex Jacobian determinant needs to be calculated to approximate the projection transformation[6]. To address this difficulty, we propose a decomposition of the terms in Eq. 7, which yields more physically interpretable components:
  • velocity of mean $\mathcal{V}_{3D}=\frac{\mathbf{V}}{\mathbf{W}}$, representing the rate of change of the 3D mean over time.
  • 3D Gaussian representation, where $\mathbf{U}-\frac{\mathbf{V}\mathbf{V}^{T}}{\mathbf{W}}$ serves as covariance matrix $\Sigma_{3D}$, and $(\mu_{x},\mu_{y},\mu_{z})^{T}$ serves as the mean vector $\mu_{3d}$
  • Temporal scaling $s_{t}=\sqrt{\mathbf{W}}$ is derived from the temporal variance term $\mathbf{W}$ in the 4D covariance matrix $\Sigma_{4D}$, which controls the variance along the time dimension and ensures it remains positive.
A detailed proof demonstrating the equivalence of these two representations is provided in the supplementary material. A key advantage of our representation is that the time-dependent variables (velocity of mean $\mathcal{V}_{3D}$, temporal scaling $s_{t}$) are represented as vectors or scalars, which can be quickly and easily projected.
Therefore, our Disentangled4DGS can be formulated as:
@@
where $\mathbf{s}_{4D}$ is the scaling vector, $q$ represents the quaternion, and $\mathcal{V}_{3D}$ denotes the velocity of mean. Our representation effectively disentangles temporal deformations, enabling for efficient and fast projection into the ray space [6]. Compared to traditional methods requiring 16 floating-point numbers, our representation only requires 15, thereby improving storage requirements by at least 4.5%.
III-B2 The Differentiable ”Projection-First” Splatting for 4D Gaussian
To enable differentiable rendering of dynamic Gaussians, we project each 4D Gaussian from world space to camera ray space. Specifically, we compute the mean, velocity, Jacobian, and covariance in ray space. First,the mean of a 4D Gaussian, denoted as $\mu_{4d}=(\mu_{3d},t)$, is tranformed from world space to camera space before projection, which can be expressed as:
$P_{3D}=\varphi(\mu_{3D}),P_{t}=\mu_{t},$ (9)